shyshy:

what is (7 2/3 + 4 3/4) - (6 5/6 + 2 3/4)=

itsme:

@supie

shyshy:

@supie plz answer

supie:

\[72/3+43/4−(65/6+23/4)=\]\[24+43/4−(65/6+23/4)=\]\[139/4−(65/6+23/4)=\]\[=139/4−199/12=?\]

shyshy:

@supie thx

supie:

np!

supie:

You can close this @shyshy.